# Theoretical study of orbital torque: Dependence on ferromagnet species and nonmagnetic layer thickness

### v1

@@ III.1 Phenomenological mechanism
Before presenting the numerical results, we first discuss the phenomenology of the current-induced torques in NM/FM bilayers. It is important to note that our calculations yield the total current-induced torque acting on the magnetization, rather than the OT contribution alone. While the damping-like torque evaluated by Eq. (5) originates from intrinsic mechanisms, multiple contributions may coexist, including OT, SOT, and self-induced torque within the FM layer. Go et al. (2020) For Ti/FM and Cu/FM bilayers, however, the SOT contribution can be safely neglected since both Ti and Cu exhibit negligible SHEs.Jo, Go, and Lee (2018); Salemi and Oppeneer (2022); Go et al. (2024) Therefore, our calculations mainly capture the OT or self-induced torque components. Distinguishing these contributions unambiguously is generally nontrivial, both theoretically and experimentally. Below, we discuss the physical mechanisms underlying these two torques and their expected dependences on the FM species and NM thickness.
@@
Meanwhile, the orbital current may originate either from the bulk or from a confined interfacial region. If it is generated only near the interface, its contribution to the OT should be largely independent of NM thickness. In contrast, if the orbital current is generated in the NM bulk, the transferred OAM, and thus the OT, should increase with increasing NM thickness, up to a characteristic length scale associated with OAM generation in the NM layer. This length scale is not necessarily identical to the orbital diffusion length. Considering that both the intrinsic OHE Go et al. (2018) and the damping-like torque described by Eq. (5) arise from interband coherence in response to the electric field, the thickness dependence may also reflect an interband coherence length. The relevant characteristic length may therefore vary with the regime of OAM transport, which remains under active debate. Rang and Kelly (2024); Sohn, Lee, and Lee (2024); Tang and Bauer (2024); Barbosa, Lee, and Rappoport (2025); Valet et al. (2025)
@@
The imbalance in spin accumulation can be induced by two different mechanisms. The first is an interfacial contribution, where inequivalent interfaces lead to different local spin accumulations at the top and bottom surfaces of the FM, irrespective of the bulk NM region. Such asymmetry can result from variations in interfacial spin scattering at the NM/FM interface, Amin, Haney, and Stiles (2020) which are typically material-specific. The resulting torque is expected to be largely independent of the NM thickness. The second is a bulk contribution, where the NM acts as a spin sink. When the spin current generated within the FM layer is injected into the NM layer, it experiences relaxation via the SOC of the NM bulk. The greater the spin absorption by the NM, the larger the resulting torque. Consequently, this contribution increases with NM thickness, where the relevant characteristic length is determined by the spin relaxation length of the NM. However, the spin relaxation length is typically significantly long in light NMs due to weak SOC, which implies that they are inefficient spin sinks. Kim, Park, and Lee (2024) If the NM layer is thinner than its spin diffusion length, the spin current will be only weakly absorbed. Therefore, in nanoscale NM/FM bilayers, the self-induced SOT is expected to be much weaker for light NMs than for heavy NMs such as Ta. Liu and Zhu (2025)
@@ III.2 Current-induced torques in Ti/FM and Cu/FM
Figure 3(a) presents the damping-like component of the torkance $t_{yx}=t_{yx}^{\mathrm{inter}}$ evaluated from Eq. (5) for Ti/FM and Cu/FM bilayers, where FM is either Co or Ni. The layer numbers are set to $N_{\mathrm{NM}}=20$ and $N_{\mathrm{FM}}=10$. All systems exhibit a negative $t_{yx}$, corresponding to the sign of the current-induced torque expected for NMs with positive spin or orbital Hall conductivity.Freimuth, Blügel, and Mokrousov (2014); Go et al. (2020) The sign of $t_{yx}$ is thus consistent with the positive orbital Hall conductivities of Ti and Cu reported in earlier theoretical works,Jo, Go, and Lee (2018); Salemi and Oppeneer (2022) although a more recent study predicted a negative orbital Hall conductivity for Cu. Go et al. (2024) However, caution is warranted, as the theoretical evaluation of orbital Hall conductivity is not yet fully established. The conventional definition of orbital Hall conductivityTanaka et al. (2008); Kontani et al. (2009); Go et al. (2018); Jo, Go, and Lee (2018); Salemi and Oppeneer (2022) is based on the conventional orbital current operator $\hat{J}_{i}^{L_{j}}=\frac{1}{2}(\hat{v}_{i}\hat{L}_{j}+\hat{L}_{j}\hat{v}_{i})$. This definition often provides qualitatively reasonable agreement with experiments, Costa et al. (2025); Santos et al. (2025) but its calculated value varies considerably across different theoretical approaches. Salemi and Oppeneer (2022); Pezo, García Ovalle, and Manchon (2022); Go et al. (2024); Rang and Kelly (2025) Moreover, the definition of $\hat{J}_{i}^{L_{j}}$ itself possesses an inherent limitation stemming from the non-conservation of OAM in crystals. Choi et al. (2023); Atencia, Arovas, and Culcer (2024); Sun et al. (2025) The orbital current undergoes a significant change due to sources such as the crystal field. Establishing a link between the conventional orbital current and experimentally measurable quantities thus requires accounting for the non-conserved part of OAM. Our approach circumvents these issues by directly computing the well-defined observable, the torque exerted on the magnetization, without explicitly invoking OAM.
Figure 3: (a) Torkance $t_{yx}$ in Ti/FM and Cu/FM bilayers (FM = Co, Ni) with $N_{\mathrm{NM}}=20$ and $N_{\mathrm{FM}}=10$. Chemical potential dependence of $t_{yx}$ in (b) Ti/FM and (c) Cu/FM bilayers.
@@
For Ti/FM bilayers, we obtain $t_{yx}=-0.16\ ea_{0}$ for Ti/Co and $t_{yx}=-0.31\ ea_{0}$ for Ti/Ni. Both values are notably large, comparable to those calculated for conventional SOT systems employing heavy NMs such as Pt and W. Freimuth, Blügel, and Mokrousov (2014); Go et al. (2020) When Ti is replaced by Cu, the torque magnitude is reduced but remains finite: $t_{yx}=-0.088\ ea_{0}$ for Cu/Co and $t_{yx}=-0.053\ ea_{0}$ for Cu/Ni. This indicates that, even without oxidation, Cu can intrinsically generate a measurable torque. The result aligns with theoretical predictions that Cu hosts a finite OHE despite its predominantly $s$-orbital character, Jo, Go, and Lee (2018); Salemi and Oppeneer (2022); Go et al. (2024) and with experimental reports of OHE in Cu, Rothschild et al. (2022); Ben Tal et al. (2025) though a quantitative comparison remains difficult.
@@
A key finding of this work is that the FM dependence of the OT in NM/FM bilayers varies depending on the NM species. For Ti/FM, Ni yields a larger torque than Co, which is consistent with experimentally observed large torques in Ti/Ni devices Hayashi et al. (2023); Choi et al. (2023); Hayashi et al. (2025) and with the established understanding that Ni’s stronger SOC enables more efficient orbital-to-spin conversion. This trend persists across a broad chemical potential range near the Fermi level, as shown in Fig. 3(b). Within the OT mechanism, the sign reversal of $t_{yx}$ with decreasing energy reflects the energy-dependent spin-orbit correlation of the FM. Go and Lee (2020); Lee et al. (2021a) The peak near 0.35-0.35 eV for Ti/Co is slightly shifted from the peak near 0.6-0.6 eV for Ti/Ni because Ni has one additional $d$ electron. Hence, the results for Ti can be reasonably explained in terms of the FM’s spin-orbit correlation.
@@
The disorder effect is modeled by introducing a constant broadening $\Gamma$ in Eq. (5). Figure 4(a) shows that in Ti/FM systems, $t_{yx}$ decreases monotonically as $\Gamma$ increases, while the relative FM dependence remains robust. The enhancement of torque when using Ni instead of Co is roughly twofold at most, consistent with its slightly stronger SOC of Ni. Figure 4(b) displays the $\Gamma$ dependence for Cu/FM systems. Interestingly, Cu/Ni exhibits a non-monotonic dependence, attributable to the sharp energy dependence of $t_{yx}$ near the Fermi level [see Fig. 3(c)]. As a result, Co yields a larger torque in the clean limit, whereas Ni produces a larger torque in the intermediate disorder regime. These results demonstrate that even for a fixed NM, the FM dependence of OT can vary, highlighting that there is no universal rule for selecting an optimal FM. Instead, the dependence varies strongly with the specific NM-FM combination and extrinsic factors such as disorder.
@@ III.3 Dependence on NM thickness: bulk origin of OT
Figures 5(a)-5(d) show $t_{yx}$ as a function of the broadening energy $\Gamma$ for various NM layer numbers $N_{\mathrm{NM}}$, while $N_{\mathrm{FM}}=10$ is fixed. For $N_{\mathrm{NM}}=1$, the NM bulk contribution is expected to be negligible, so the torque mainly originates from the NM/FM interface. As $N_{\mathrm{NM}}$ increases, the bulk contribution becomes more pronounced, while the interfacial contribution would remain nearly unchanged. Comparing the results for $N_{\mathrm{Ti}}=1$ and $N_{\mathrm{Ti}}=40$ [Figs. 5(a) and  5(b)], we find that the bulk Ti contribution dominates once the Ti layer becomes sufficiently thick for both Ti/Co and Ti/Ni. In particular, the thickness dependence becomes more prominent as $\Gamma$ decreases, while it weakens for $\Gamma\gtrsim 100$ meV.
Noticeable thickness dependences of $t_{yx}$ are also observed in Cu/FM bilayers [Figs. 5(c) and  5(d)]. Interestingly, for Cu/Ni, the variation of $t_{yx}$ with $N_{\mathrm{Cu}}$ reverses sign between clean and disordered regimes. For $\Gamma\lesssim 30$ meV, $t_{yx}$ increases in the positive direction with $N_{\mathrm{Cu}}$, implying that the Cu bulk contribution in this regime has the opposite sign to the OT expected for a positive orbital Hall conductivity. Consequently, the total torque becomes positive in the thick limit, unlike in other systems. This sign reversal recalls the ambiguity in the theoretically predicted sign of the orbital Hall conductivity of Cu. In addition, as discussed in Sec. III.2, the conventional definition of the orbital Hall conductivity does not account for the non-conservation of OAM in crystals, an effect that is implicitly taken into account in the torque calculation. Go et al. (2020) In this regard, our results reveal a possible discrepancy between the theoretically calculated bulk orbital Hall conductivity and the effective orbital Hall conductivity measured in OT experiments.
@@
To more clearly visualize the thickness dependence, we plot $t_{yx}$ as a function of $N_{\mathrm{NM}}$ for two selected broadening energies, $\Gamma=10$ meV and $\Gamma=100$ meV, corresponding to the clean and disordered regimes, respectively [Figs. 5(e)-5(h)]. The torque is strongly suppressed for $N_{\mathrm{NM}}=1$, indicating a minor interfacial contribution in all cases. This aligns with experiments that report negligible damping-like torque in NM/Py devices with ultrathin ($\sim 0.5$ nm) Ti and Cu layers. Greening et al. (2020) For Ti/FM [Fig. 5(f)], the torque saturates rapidly, which seems to agree with recent first-principles results showing a fast decay of orbital currents within a few atomic layers. Rang and Kelly (2024) However, we note that Ref. Rang and Kelly (2024) investigated orbital currents injected from the FM, which differs from the OT mechanism considered here. Indeed, Fig. 5(e) shows that the torque does not fully saturate even for $N_{\mathrm{Ti}}=40$ ($\approx 9.6$ nm). Such a long characteristic length—well beyond atomic scales—shows qualitative agreement with experimental observations of long-range OT in Ti/FM devices. Choi et al. (2023); Hayashi et al. (2023); Xu et al. (2025a); Hayashi et al. (2025); Xu et al. (2025b)
@@
Finally, we emphasize that the NM bulk contribution mainly originates from the OT. Although both OT and self-induced SOT can, in principle, depend on NM thickness (see Sec. III.1), the latter contribution is likely minor for two reasons. First, first-principles calculations show that the spin Hall conductivity of ferromagnetic Ni is significantly large, comparable to that of Pt, whereas that of ferromagnetic Co is negligible. Go et al. (2024) In our calculations, however, the torques in NM/Co and NM/Ni are of the same order of magnitude. Second, the length scale of a self-induced SOT is governed by the spin diffusion length of the NM, which can reach several hundred nanometers in light NMs such as Cu. Yakata et al. (2006) By contrast, the thickness dependence observed here extends only up to about 10 nm, suggesting a much shorter relevant length scale. Meanwhile, we note that the possibility of a self-induced OT cannot be entirely ruled out. The orbital Hall conductivities for Co and Ni are comparable, Salemi and Oppeneer (2022); Go et al. (2024) and an orbital diffusion length on the order of a few nanometers is plausible. A clear identification and separation of such effects are left for future work.

### v2

@@ III.1 Phenomenological mechanism
Before presenting the numerical results, we first discuss the phenomenology of the current-induced torques in NM/FM bilayers. It is important to note that our calculations yield the total current-induced torque acting on the magnetization, rather than the OT contribution alone. While the damping-like torque evaluated by Eq. (8) originates from intrinsic mechanisms, multiple contributions may coexist, including OT, SOT, and self-induced torque within the FM layer. Go et al. (2020) For Ti/FM and Cu/FM bilayers, however, the SOT contribution can be safely neglected since both Ti and Cu exhibit negligible SHEs.Jo et al. (2018); Salemi and Oppeneer (2022); Go et al. (2024) Therefore, our calculations mainly capture the OT or self-induced torque components. Distinguishing these contributions unambiguously is generally nontrivial, both theoretically and experimentally. Below, we discuss the physical mechanisms underlying these two torques and their expected dependences on the FM species and NM thickness.
@@
Meanwhile, the orbital current may originate either from the bulk or from a confined interfacial region. If it is generated only near the interface, its contribution to the OT should be largely independent of NM thickness. In contrast, if the orbital current is generated in the NM bulk, the transferred OAM, and thus the OT, should increase with increasing NM thickness, up to a characteristic length scale associated with OAM generation in the NM layer. This length scale is not necessarily identical to the orbital diffusion length. Considering that both the intrinsic OHE Go et al. (2018) and the damping-like torque described by Eq. (8) arise from interband coherence in response to the electric field, the thickness dependence may also reflect an interband coherence length. The relevant characteristic length may therefore vary with the regime of OAM transport, which remains under active debate. Rang and Kelly (2024); Sohn et al. (2024); Tang and Bauer (2024); Barbosa et al. (2025); Valet et al. (2025)
@@
The imbalance in spin accumulation can be induced by two different mechanisms. The first is an interfacial contribution, where inequivalent interfaces lead to different local spin accumulations at the top and bottom surfaces of the FM, irrespective of the bulk NM region. Such asymmetry can result from variations in interfacial spin scattering at the NM/FM interface, Amin et al. (2020) which are typically material-specific. The resulting torque is expected to be largely independent of the NM thickness. The second is a bulk contribution, where the NM acts as a spin sink. When the spin current generated within the FM layer is injected into the NM layer, it experiences relaxation via the SOC of the NM bulk. The greater the spin absorption by the NM, the larger the resulting torque. Consequently, this contribution increases with NM thickness, where the relevant characteristic length is determined by the spin relaxation length of the NM. However, the spin relaxation length is typically significantly long in light NMs due to weak SOC, which implies that they are inefficient spin sinks. Kim et al. (2024) If the NM layer is thinner than its spin diffusion length, the spin current will be only weakly absorbed. Therefore, in nanoscale NM/FM bilayers, the self-induced SOT is expected to be much weaker for light NMs than for heavy NMs such as Ta. Liu and Zhu (2025)
@@ III.2 Current-induced torques in Ti/FM and Cu/FM
Figure 3(a) presents the damping-like component of the torkance $t_{yx}=t_{yx}^{\mathrm{inter}}$ evaluated from Eq. (8) for Ti/FM and Cu/FM bilayers, where FM is either Co or Ni. The layer numbers are set to $N_{\mathrm{NM}}=20$ and $N_{\mathrm{FM}}=10$. All systems exhibit a negative $t_{yx}$, corresponding to the sign of the current-induced torque expected for NMs with positive spin or orbital Hall conductivity.Freimuth et al. (2014); Go et al. (2020) The sign of $t_{yx}$ is thus consistent with the positive orbital Hall conductivities of Ti and Cu reported in earlier theoretical works,Jo et al. (2018); Salemi and Oppeneer (2022) although a more recent study predicted a negative orbital Hall conductivity for Cu. Go et al. (2024) However, caution is warranted, as the theoretical evaluation of orbital Hall conductivity is not yet fully established. The conventional definition of orbital Hall conductivityTanaka et al. (2008); Kontani et al. (2009); Go et al. (2018); Jo et al. (2018); Salemi and Oppeneer (2022) is based on the conventional orbital current operator $\hat{Q}_{i}^{L_{j}}=\frac{1}{2}(\hat{v}_{i}\hat{L}_{j}+\hat{L}_{j}\hat{v}_{i})$. This definition often provides qualitatively reasonable agreement with experiments, Costa et al. (2025); Santos et al. (2025) but its calculated value varies considerably across different theoretical approaches. Salemi and Oppeneer (2022); Pezo et al. (2022); Go et al. (2024); Rang and Kelly (2025) Moreover, the definition of $\hat{Q}_{i}^{L_{j}}$ itself possesses an inherent limitation stemming from the non-conservation of OAM in crystals, Choi et al. (2023); Atencia et al. (2024); Sun et al. (2025) as discussed in Sec. II.2. The orbital current undergoes a significant change due to sources such as the crystal field. Establishing a link between the conventional orbital current and experimentally measurable quantities thus requires accounting for the non-conserved part of OAM [see, e.g., Eq. (3)]. Our approach circumvents these issues by directly computing the well-defined observable, the torque exerted on the magnetization, without explicitly invoking OAM.
Figure 3: (a) Torkance $t_{yx}$ in Ti/FM and Cu/FM bilayers (FM = Co, Ni) with $N_{\mathrm{NM}}=20$ and $N_{\mathrm{FM}}=10$. Chemical potential dependence of $t_{yx}$ in (b) Ti/FM and (c) Cu/FM bilayers. Dotted lines represent the torque in hypothetical systems where the chemical potential of the NM layer is shifted so that the OHE is suppressed. The energy broadening is assumed to be $\Gamma=25$ meV.
@@
For Ti/FM bilayers, we obtain $t_{yx}=-0.18\ ea_{0}$ for Ti/Co and $t_{yx}=-0.33\ ea_{0}$ for Ti/Ni. Both values are notably large, comparable to those calculated for conventional SOT systems employing heavy NMs such as Pt and W. Freimuth et al. (2014); Go et al. (2020) When Ti is replaced by Cu, the torque magnitude is reduced but remains finite: $t_{yx}=-0.11\ ea_{0}$ for Cu/Co and $t_{yx}=-0.047\ ea_{0}$ for Cu/Ni. This indicates that, even without oxidation, Cu can intrinsically generate a measurable torque. The result aligns with theoretical predictions that Cu hosts a finite OHE despite its predominantly $s$-orbital character, Jo et al. (2018); Salemi and Oppeneer (2022); Go et al. (2024) and with experimental reports of OHE in Cu, Rothschild et al. (2022); Ben Tal et al. (2025) though a quantitative comparison remains difficult.
@@
A key finding of this work is that the FM dependence of the OT in NM/FM bilayers varies depending on the NM species. For Ti/FM, Ni yields a larger torque than Co, which is consistent with experimentally observed large torques in Ti/Ni devices Hayashi et al. (2023); Choi et al. (2023); Hayashi et al. (2025) and with the established understanding that Ni’s stronger SOC enables more efficient orbital-to-spin conversion. This trend persists across a broad chemical potential range near the Fermi level, as shown in Fig. 3(b). Although it is impractical to drastically tune the chemical potential in metallic systems, examining the energy dependence can still provide useful insights. Within the OT mechanism, the sign reversal of $t_{yx}$ with decreasing energy reflects the energy-dependent spin-orbit correlation of the FM. Go and Lee (2020); Lee et al. (2021a) The peak near 0.35-0.35 eV for Ti/Co is slightly shifted from the peak near 0.6-0.6 eV for Ti/Ni because Ni has one additional $d$ electron. Hence, the results for Ti can be reasonably explained in terms of the FM’s spin-orbit correlation.
@@
The disorder effect is modeled by introducing a constant broadening $\Gamma$ in Eq. (8). In this formalism, $\Gamma$ is associated with a finite quasiparticle lifetime, $\tau=\hbar/\Gamma$, which phenomenologically accounts for scattering in the sample. Increasing $\Gamma$ shortens $\tau$, thereby moving the system from a cleaner to a more disordered regime and making interband coherence more easily destroyed. Figure 4(a) shows that in Ti/FM systems, $|t_{yx}|$ decreases monotonically as $\Gamma$ increases, while the relative FM dependence remains robust. The enhancement of torque when using Ni instead of Co is roughly twofold at most, consistent with its slightly stronger SOC of Ni. Figure 4(b) displays the $\Gamma$ dependence for Cu/FM systems. Interestingly, Cu/Ni exhibits a non-monotonic dependence, attributable to the sharp energy dependence of $t_{yx}$ near the Fermi level [see Fig. 3(c)]. As a result, Co yields a larger torque in the clean limit, whereas Ni produces a larger torque in the intermediate disorder regime. These results demonstrate that even for a fixed NM, the FM dependence of OT can vary, highlighting that there is no universal rule for selecting an optimal FM. Instead, the dependence varies strongly with the specific NM-FM combination and extrinsic factors such as disorder.
Figure 5: (a) Total DOS of bulk Ti, Cu, Co, and Ni. (b)-(d) Projected DOS for the (b) $s$, (c) $p$, and (d) $d$ orbitals. The Fermi level is set to 0 eV.
To gain insight into the effect of NM-FM matching on the OT, we compare the bulk density of states (DOS) of the constituent materials [Figs. 5(a)-5(d)]. The DOS of Ti shows no notable difference in its overlap with the DOS of Co versus Ni. By contrast, the total DOS of Cu overlaps more strongly with that of the spin-majority bands of Co than with that of Ni. This feature mainly originates from the $d$-orbital states [Fig. 5(d)], which are expected to dominate the orbital response. Because the damping-like torque is obtained from a sum over occupied states [Eq. (8)], not only the matching at the Fermi energy but also that over an energy window below it can influence the torque. Therefore, this result suggests that the FM dependence of OT can be affected by the specific NM-FM material combination in addition to the intrinsic properties of the FM layer itself.
We further investigate the possibility of a self-induced SOT from the FM layer. To estimate this contribution, we consider hypothetical NM systems in which the $d$-orbital character is suppressed and the OHE is therefore strongly reduced. Specifically, we shift the chemical potential of Ti and Cu downward by 4 eV and 7 eV, respectively, and denote these systems as Ti and Cu. Figures 5(b)-5(d) indicate that the valence electrons in Ti and Cu are dominated by $s$-orbital character. In Ti/FM and Cu/FM bilayers, the OHE contribution from the NM is thus expected to be negligible, while the self-induced SOT from the FM may remain. The calculated current-induced torques in these systems are shown by the dotted lines in Figs. 3(b) and 3(c), respectively. We find that both the Ti/FM and Cu/FM bilayers exhibit negligible torque at the Fermi energy, indicating that the self-induced SOT is not a dominant contribution to the torque.
So far we have focused on the exchange torque [Eq. (4)], which represents the transfer of angular momentum from conduction electrons to the local spin moment of the FM layer. Meanwhile, another important issue is how much the generated OAM is absorbed into the lattice. The OAM is normally expected to be largely quenched in crystals. To investigate this, we recall the continuity equation for angular momentum discussed in Sec. II.2. Integrating spatially Eq. (3) over the entire sample, the flux of the angular momentum current vanishes; thus, we obtain the following relation in terms of spatially averaged values:
$\mathbf{T}^{\mathrm{XC}}=-\frac{\partial\mathbf{J}}{\partial t}+\mathbf{T}^{\mathrm{CF}}.$ (9)
This equation states that the exchange torque is balanced by the net accumulation rate and the crystal field torque. Note that this relation holds only at the global level, not locally. In the linear regime, the current-induced responses can be written as $T_{y}^{\mathrm{XC/CF}}=t_{yx}^{\mathrm{XC/CF}}E_{x}$ and $J_{y}=\chi_{yx}E_{x}$, which approximately yields
$t_{yx}^{\mathrm{XC}}\approx-\frac{\chi_{yx}}{\tau}+t_{yx}^{\mathrm{CF}},$ (10)
where $\tau=\hbar/\Gamma$ is the lifetime. While $t_{yx}^{\mathrm{XC}}$ is evaluated by Eqs. (7) and (8), $\chi_{yx}$ and $t_{yx}^{\mathrm{CF}}$ are obtained by replacing $\hat{T}_{y}^{\mathrm{XC}}$ with $\hat{J}_{y}$ and $\hat{T}_{y}^{\mathrm{CF}}$, respectively. The calculated values are shown in Figs. 6(a)-6(d). The results indicate that $t_{yx}^{\mathrm{CF}}$ is much larger than $t_{yx}^{\mathrm{XC}}$ and very close to $\chi_{yx}/\tau$ in general. This implies that a large portion of angular momentum is absorbed into the lattice, as typically expected, and only a relatively smaller part is transferred to the magnetization. It is worth noting that $t_{yx}^{\mathrm{XC}}$ is nevertheless still sizable. This also suggests the possibility of enhancing the OT by suppressing the crystal field torque, which may be of potential interest for orbitronic engineering.
Figure 6: Calculated values for each term appearing in Eq. (10) for (a) Ti/Co, (b) Ti/Ni, (c) Cu/Co, and (d) Cu/Ni bilayers. The energy broadening is assumed to be $\Gamma=25$ meV.
@@ III.3 Dependence on NM thickness: bulk origin of OT
Figures 7(a)-7(d) show $t_{yx}$ as a function of the broadening energy $\Gamma$ for various NM layer numbers $N_{\mathrm{NM}}$, while $N_{\mathrm{FM}}=10$ is fixed. For $N_{\mathrm{NM}}=1$, the NM bulk contribution is expected to be negligible, so the torque mainly originates from the NM/FM interface. As $N_{\mathrm{NM}}$ increases, the bulk contribution becomes more pronounced, while the interfacial contribution would remain nearly unchanged. Comparing the results for $N_{\mathrm{Ti}}=1$ and $N_{\mathrm{Ti}}=40$ [Figs. 7(a) and  7(b)], we find that the bulk Ti contribution dominates once the Ti layer becomes sufficiently thick for both Ti/Co and Ti/Ni. In particular, the thickness dependence becomes more prominent as $\Gamma$ decreases, while it weakens for $\Gamma\gtrsim 100$ meV.
Noticeable thickness dependences of $t_{yx}$ are also observed in Cu/FM bilayers [Figs. 7(c) and  7(d)]. Interestingly, for Cu/Ni, the variation of $t_{yx}$ with $N_{\mathrm{Cu}}$ reverses sign between clean and disordered regimes. For $\Gamma\lesssim 30$ meV, $t_{yx}$ increases in the positive direction with $N_{\mathrm{Cu}}$, implying that the Cu bulk contribution in this regime has the opposite sign to the OT expected for a positive orbital Hall conductivity. Consequently, the total torque becomes positive in the thick limit, unlike in other systems. This sign reversal recalls the ambiguity in the theoretically predicted sign of the orbital Hall conductivity of Cu. In addition, as discussed in Secs. II.2 and III.2, the conventional definition of the orbital current is not directly related to the torque acting on the magnetization. The conventional orbital Hall conductivity can serve as a good indicator only when the crystal field torque appearing in Eq. (3) or Eq. (9) is sufficiently small. In this regard, our results reveal the existence of a discrepancy between the theoretically calculated bulk orbital Hall conductivity and the effective orbital Hall conductivity measured in OT experiments.
@@
To more clearly visualize the thickness dependence, we plot $t_{yx}$ as a function of $N_{\mathrm{NM}}$ for two selected broadening energies, $\Gamma=10$ meV and $\Gamma=100$ meV, corresponding to the clean and disordered regimes, respectively [Figs. 7(e)-7(h)]. The torque is strongly suppressed for $N_{\mathrm{NM}}=1$, indicating a minor interfacial contribution in all cases. This aligns with experiments that report negligible damping-like torque in NM/Py devices with ultrathin ($\sim 0.5$ nm) Ti and Cu layers. Greening et al. (2020) For Ti/FM [Fig. 7(f)], the torque saturates rapidly, which seems to agree with recent first-principles results showing a fast decay of orbital currents within a few atomic layers. Rang and Kelly (2024) However, we note that Ref. Rang and Kelly (2024) investigated orbital currents injected from the FM, which differs from the OT mechanism considered here. Indeed, Fig. 7(e) shows that the torque does not fully saturate even for $N_{\mathrm{Ti}}=40$ ($\approx 9.6$ nm). Such a long characteristic length—well beyond atomic scales—shows qualitative agreement with experimental observations of long-range OT in Ti/FM devices. Choi et al. (2023); Hayashi et al. (2023); Xu et al. (2025b); Hayashi et al. (2025); Xu et al. (2025a)
@@
Finally, we emphasize that the NM bulk contribution mainly originates from the OT. Although both OT and self-induced SOT can, in principle, depend on NM thickness (see Sec. III.1), the latter contribution is likely minor for the following reasons. First, when the OHE in the NM layer is suppressed, the torque becomes negligible, as shown in Figs. 3(b) and 3(c). Second, first-principles calculations show that the spin Hall conductivity of ferromagnetic Ni is significantly large, comparable to that of Pt, whereas that of ferromagnetic Co is negligible. Go et al. (2024) In our calculations, however, the torques in NM/Co and NM/Ni are of the same order of magnitude. Lastly, the length scale of a self-induced SOT is governed by the spin diffusion length of the NM, which can reach several hundred nanometers in light NMs such as Cu. Yakata et al. (2006) By contrast, the thickness dependence observed here extends only up to about 10 nm, suggesting a much shorter relevant length scale. Meanwhile, we note that the possibility of a self-induced OT cannot be entirely ruled out. The orbital Hall conductivities for Co and Ni are comparable, Salemi and Oppeneer (2022); Go et al. (2024) and an orbital diffusion length on the order of a few nanometers is plausible. A clear identification and separation of such effects are left for future work.